An anti-air craft gun can take a maximum of 4 shots at an enemy plane moving away from it. The probabilities of hitting the plane at the first, second,…
An anti-air craft gun can take a maximum of 4 shots at an enemy plane moving away from it. The probabilities of hitting the plane at the first, second, third and fourth shot are 0.6, 0.5, 0.4 and 0.2 respectively. What is the probability that the gun hit the plane?
Show answerHide answer
Correct answer: (A) 0.904
Explanation
The chance that all four shots miss is 0.4 × 0.5 × 0.6 × 0.8 = 0.096. So P(at least one hit) = 1 − 0.096 = 0.904.